<Chapter 17 - Homework 5 or 33> Forces in a Three-Charge System What is the net force exerted by these two charges on a third charge 50.0 nC placed between i and g2 at Constants 1.070 m? Coulombs law for the magnitude of the force F botween two particles with charges Q and Q separated by a distance d is Your answer may be positive or negative, depending on the direction of the force Express your answer numerically in newtons to three significant tigures. d2 View Available Hint(s) where K-and «o = 8.854 × 10 12 C?/(N·m?) is the permittivity of free space Consider two point charges loc ated on the x axis one charge, h17.5 nC, is located at z1·1 650 m : the second charge, g2 = 34 5 nC is at the origin (a 0.0000) Force Submit incorrect; Try Again: 3 attempts remaining
Answer #1

Net Force on charge q3 will be:

Force F13 will be attractive means towards -ve x-axis

Force F23 will be repulsive which means towards -ve x-axis

So,

F3 = - F13 - F23

F3 = -k*q1*q3/r13^2 - k*q2*q3/r23^2

r13 = distance between q1 and q3 = (-1.070 - (-1.650)) = 0.58 m

r23 = distance between q2 and q3 = (-1.070 - (0)) = -1.070 m

k = 1/(4*pi*e0) = 9*10^9

So,

F3 = [-9*10^9*17.5*10^-9*50*10^-9/0.58^2] - 9*10^9*34.5*10^-9*50*10^-9/(-1.070)^2

F3 = -36.7*10^-6 N

Please Upvote.

Comment below if you have any query.